# Unified Lorentz-covariant Poisson Bracket for the Electrodynamics of a Point Particle

### v1

@@ Abstract
Using the multisymplectic Hamiltonian formalism, we propose a Poisson bracket for the electromagnetic field that, in addition to satisfying the restricted principle of relativity, reproduces well-established results from the standard formulations of field theories. To this end, it is necessary to translate the multisymplectic description of the system to the field’s momentum representation, where a bilinear and Lorentz-covariant form emerges in a straightforward manner. Consequently, we establish a description for a system consisting of an electromagnetic field and a point particle interacting with each other that is based on a single Lorentz-covariant Poisson bracket with respect to all the canonical variables of the system. The results obtained from this proposal offer a solid foundation for the development of a Lorentz-covariant quantization process.
Keywords: Field theory, classical electrodynamics, multisymplectic Hamiltonian formalism, Lorentz-covariant Poisson bracket
@@ 1 Introduction
Since the early objections raised by Born [1], the incompatibility between the requirement of full consistency with the restricted principle of relativity and the use of canonical quantization in modern field theories has been addressed on numerous occasions without yet reaching a definitive solution. The crux of the matter lies in that canonical quantization requires a Hamiltonian (or canonical) formulation of the theories and, as a consequence, implies a primacy of time as fundamental independent variable. This incompatibility is only apparent, naturally, since quantum field theories do describe relativistic phenomena. However, the covariance of such theories must be verified at each stage of their development by means of generally laborious demonstrations.
Constructing a Lorentz-covariant bilinear form that generalizes the Poisson bracket (pb) of classical mechanics when dealing with fields is another open question in theoretical physics that is closely related to the previous issue. In this regard, the studies done by Peierls [2] and by Marsden et al. [3] are of great significance, as they clearly illustrate that there is no unique or straightforward way to construct a pb with the aforementioned characteristics. Multiple proposals for Lorentz-covariant pbs have been published over the years for both classical and quantum fields (see, for example, [2, 3, 6, 7, 8, 9, 10, 4, 5] and references therein), indicating that no general consensus exists on this matter.
@@
Hence, generalizing the Hamiltonian formalism in such a way that it can underpin a manifestly Lorentz-covariant quantization process is a topic of major interest. A strong candidate for this generalization is the multisymplectic Hamiltonian formalism (mhf), whose main feature is to manifestly preserve the covariance of the canonical description of fields through the introduction of a generalized momentum per field derivative. The seminal papers on the mhf were published by de Donder [11] and by Weyl [12] in the 1930s. However, after nearly a century of existence, the use of this formalism to construct field theories is not common practice, since its development beyond the basic definitions is not straightforward and still presents formal lacunae [13]. In particular, the mhf has as an important drawback that it allows to define more than one pb [13, 14, 15, 9, 16, 17]. Nonetheless, it has recently come to our attention that it is possible to construct a Lorentz-covariant pb in a straightforward manner for the electromagnetic field by translating the multisymplectic description of the field to its momentum representation [18]. Exploring this possibility is the main subject of the present work.
@@
The paper is organized as follows. In Sect. 2, after briefly reviewing the fundamentals of describing a system composed of an electromagnetic field and a single point particle using the mhf, we translate the generated description to the momentum representation. The central part of the paper is contained in Sect. 3 and deals with the proposal of a pb with respect to the canonical variables of the field that is capable not only of satisfying the restricted principle of relativity but also of reproducing well-established results from the conventional formulations of field theories. Subsequently, in Sect. 4, the non-covariant form of the proposal is analyzed and compared with the non-covariant pb that is associated with the electromagnetic field. Lastly, we outline a description of the complete system on the basis of the found pb and briefly discuss its main implications in Sect. 5.
@@ 2 Multisymplectic Hamiltonian Description
We start by properly setting up the system to be investigated as an electromagnetic field and a single point particle interacting with each other in a four-dimensional Minkowski space-time, with metric tensor $\eta_{\mu\nu}$ and signature (+,,,)(+,-,-,-). The action of such a system can be stated as
@@
where $\mathcal{L}_{\mathrm{field}}$, $\mathcal{L}_{\mathrm{part}}$, and $\mathcal{L}_{\mathrm{int}}$ are the Lagrangian densities associated with the free propagation of the field, the free motion of the particle, and the particle++field interaction, respectively. Thus, we have that
$\mathcal{L}_{\mathrm{field}}=-\frac{1}{8\pi{c}}\partial_{\mu}A_{\nu}\partial^{\mu}A^{\nu},$ (2)
$\mathcal{L}_{\mathrm{part}}=-\frac{m_{0}c^{2}}{2}\int{d}\tau\hskip 1.42262pt\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (3)
@@
with $A_{\mu}$ the field’s vector potential and $m_{0}$, $e$, $u_{\mu}$, and $\tau$ the rest mass, electric charge, position, and proper time of the particle, respectively [^2] The Lagrangian density in Eq. (2) does not correspond to the gauge-invariant form commonly used in electrodynamics that is given in terms of the electromagnetic tensor. However, both Lagrangian densities differ only by a quantity that reduces to zero when integrated over the whole space-time and when considering the Lorentz condition, $\partial_{\mu}A^{\mu}=0$, so that the selection of one option or the other does not modify the action of the system..
Thereupon, the mhf implies a generalized conjugate momentum for $A_{\mu}$ given by
$\theta_{\mu\nu}=-\frac{1}{4\pi{c}}\partial_{\mu}A_{\nu},$ (5)
while the application of the usual Hamiltonian formalism results for the mechanical subsystem in the well-known canonical momentum
$p_{\mu}=-m_{0}\dot{u}_{\mu}-\frac{e}{c}A_{\mu}(u_{\alpha}).$ (6)
Consequently, we obtain that the de Donder-Weyl function, i.e., the Lorentz-covariant Hamiltonian density, for the system is
$\mathcal{H}=-2\pi{c}\hskip 0.7113pt\theta_{\mu\nu}\theta^{\mu\nu}-\frac{1}{2m_{0}}\int{d}\tau\left[p_{\mu}+\frac{eA_{\mu}}{c}\right]\left[p^{\mu}+\frac{eA^{\mu}}{c}\right]\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (7)
which leads to two pairs of Lorentz-covariant first-order Hamilton’s equations: one pair for the field’s variables and the other for those of the point particle. The set of these equations consistently reproduces the manifestly Lorentz-covariant forms of Maxwell’s equations and the particle’s equation of motion.
@@
$A_{\nu}=\frac{1}{8\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha}){\mathcal{A}}_{\nu}e^{-{i}k_{\mu}x^{\mu}}+\mathrm{c.c.},$ (8)
where the Dirac $\delta$-function $\delta(k_{\mu}k^{\mu})$ and the Heaviside step function $\Theta(k_{0})$ ensure that the dispersion relation $k_{0}=|\mathbf{k}|$ is satisfied, and ${\mathcal{A}}_{\nu}$ is a vector amplitude given by
${\mathcal{A}}_{\nu}={\mathcal{A}}^{\mathrm{free}}_{\nu}+4\pi{i}e\int{d}\tau\hskip 1.42262pt\dot{u}_{\nu}\Theta(x_{0}-u_{0})e^{ik_{\mu}u^{\mu}},$ (9)
with $\mathcal{A}_{\mu}^{\mathrm{free}}$ a vector amplitude depending solely on $k_{\mu}$. Having expressed the solution of the field equation in this manner, one can easily observe that the presence of the particle results for the field in a superposition of plane waves with constant amplitudes due to its free propagation and of waves with variable amplitudes whose origin can be shown to be the radiation reaction of the particle (for a detailed proof, see [19]).
Then, recalling that the description of the field provided by the mhf in the momentum representation has been shown to be equivalent to its Lagrangian counterpart [18], we have that Eq. (7) can be rewritten as
@@
with $\delta_{k}$ a $\delta$-function such that $\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\delta_{k}=8\pi^{3}$, is the image of the system’s de Donder-Weyl function in the field’s momentum representation. This last function has the same information about the system as $\mathcal{H}$ and therefore constitutes the basis for this alternative description. Furthermore, the structure of Eq. (11) reveals that the de Donder-Weyl function is a Lorentz scalar proportional to the rest mass of the complete system, i.e., particle++field [^3] The origin of the proposed interpretation becomes evident when recalling that the dispersion relation for a massive field is $k_{\mu}k^{\mu}=\mu^{2}$, with $\mu$ being the mass parameter associated with the field. For purposes of comparison, examples of de Donder-Weyl functions in the momentum representation for other fields may be seen in Ref. [18]. Naturally, this identification implies a null value for the de Donder-Weyl function’s part associated with the free propagation of the electromagnetic field since the dispersion relation states in such a case that $k_{\mu}k^{\mu}=k_{0}^{2}-|\mathbf{k}|^{2}=0$.
The image of the de Donder-Weyl function in the momentum representation can equivalently be expressed in terms of the field’s canonical variables ${q}_{\mu}$ and $\pi_{\mu\nu}$ [18], defined as
${q}_{\mu}=\frac{i(\mathcal{A}_{\mu}^{\ast}-\mathcal{A}_{\mu})}{\sqrt{8\pi{c}k_{0}}},\quad\pi_{\mu\nu}=\frac{k_{\mu}(\mathcal{A}_{\nu}^{\ast}+\mathcal{A}_{\nu})}{\sqrt{8\pi{c}k_{0}}},$ (12)
and whose Hamilton’s equations are:
@@
respectively. As a consequence, the set of equations constituted by Eqs. (13) and Hamilton’s equations for $u_{\mu}$ and $p_{\mu}$ completely governs the electrodynamics of the whole system.
@@ 3 Proposal for a Lorentz-covariant Poisson Bracket
In the Hamiltonian formulation of classical mechanics, the evolution equation of any phase-space function is stated in terms of its pb with the Hamiltonian function of the mechanical system concerned, with Hamilton’s equations being special cases of such an evolution equation for the canonical positions and momenta. In particular, when considering the Hamiltonian function itself, one immediately finds that its evolution equation reduces to an equality between its total and explicit time derivatives. So, if we claim that the above description of the system is canonical, we must verify whether the de Donder-Weyl function satisfies a similar mathematical relationship between its derivatives. To this end, we observe that it can be established straightforwardly from Eqs. (10) and (13) that
@@
so that we conclude that the de Donder-Weyl function indeed satisfies an equality such as the one mentioned.
The former statement brings pbs back to the center of our discussion. As mentioned in Sect. 1, the mhf accepts multiple ways of defining such geometric objects, yet the structure of Eq. (14) strongly suggests proposing the bilinear form
$\{A,B\}=\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})V_{\mu}\left[\frac{\partial{A}}{\partial{q}^{\nu}}\frac{\partial{B}}{\partial\pi_{\mu\nu}}-\frac{\partial{B}}{\partial{q}^{\nu}}\frac{\partial{A}}{\partial\pi_{\mu\nu}}\right],$ (15)
@@
$\{q_{\mu}(x_{0},k_{\alpha}),\pi_{\lambda\nu}(x_{0},k_{\alpha}^{\prime})\}=\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})V_{\lambda}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}),$ (16)
@@
In order to get a hint on the structure of $V_{\mu}$, it is very enlightening to calculate the equal-time pb of the field and its (generalized) conjugate momentum:
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{1}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{\mu\nu}K_{\lambda}(\mathbf{x},\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (17)
@@
$K_{\mu}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{4\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})k_{0}^{-1}\left[V_{\mu}\sin(\mathbf{k}\cdot\mathbf{x})\sin(\mathbf{k}\cdot\mathbf{x}^{\prime})+\frac{k_{\alpha}V^{\alpha}}{k_{\alpha}k^{\alpha}}k_{\mu}\cos(\mathbf{k}\cdot\mathbf{x})\cos(\mathbf{k}\cdot\mathbf{x}^{\prime})\right],$ (18)
since this result is independent of the auxiliary variables used to describe the system. Thus, Eq. (17) calls for taking $V_{\mu}$ equal to $ak_{0}k_{\mu}$, with $a$ a proportionality constant, so that
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{a}{4}\left[\frac{\delta^{2}(0)}{8\pi^{3}}\right]\eta_{0\lambda}\eta_{\mu\nu}\delta^{(3)}(\mathbf{x}-\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (19)
With this selection, we easily identify the well-known result from standard field theory for the mentioned pb in the part of the latter equation that is associated with the free propagation of the electromagnetic field.

### v2

@@ Abstract
Using the multisymplectic Hamiltonian formalism, we propose a Poisson bracket for an electromagnetic radiation field that, in addition to satisfying the restricted principle of relativity, reproduces well-established results from the conventional formulations of field theories. To this end, it is necessary to translate the multisymplectic description of the system to the field’s momentum representation, where a bilinear and Lorentz-covariant form emerges in a straightforward manner. Consequently, we establish a description for a system consisting of an electromagnetic radiation field and a point particle interacting with each other that is based on a single Lorentz-covariant Poisson bracket with respect to the canonical variables of the system. The results obtained from this proposal offer a solid foundation for the development of a Lorentz-covariant quantization process.
Keywords: Canonical quantization, multisymplectic Hamiltonian formalism, Lorentz-covariant Poisson bracket
@@ 1 Introduction
Since the early objections raised by Born [1], the incompatibility between the requirement of full consistency with the restricted principle of relativity and the use of canonical quantization in modern field theories has been addressed on numerous occasions without reaching yet a definitive solution. The crux of the matter lies in the fact that canonical quantization requires a Hamiltonian (or canonical) formulation of field theories and therefore implies a primacy of time as fundamental independent variable [2, 3]. This incompatibility is not fundamental: quantum field theories constructed through canonical quantization do describe relativistic phenomena; however, their covariance must be verified at every stage of construction, typically through laborious proofs.
Alternative quantization processes that avoid privileging time and thus yield quantum field theories in a manifestly Lorentz-covariant manner have been developed over the years. The functional (or path-integral) quantization process, for example, formulates quantum field theories in terms of integrals over field configurations using the action of the system as weight function [4, 5, 6]. Unlike the Hamiltonian function, the action is a Lorentz scalar, so functional quantization preserves covariance throughout the formulation of the theories. Moreover, the Faddeev–Popov procedure, Becchi–Rouet–Stora–Tyutin quantization, and the Batalin–Vilkovisky formalism provide systematic, manifestly Lorentz-covariant methods for quantizing gauge theories [7, 8, 9, 10]. The Haag–Kastler axiomatic framework, in which fields are defined as operator-valued distributions subject to conditions that include covariance, offers, at least in principle, another Lorentz-covariant route for formulating quantum field theories [11]. Yet, despite these alternatives, constructing a fully Lorentz-covariant canonical quantization process remains an open problem in theoretical physics.
@@
The starting point for addressing this problem has historically been the development of a Lorentz-covariant generalization of the Hamiltonian formalism. In this regard, a strong candidate for such a generalization is the multisymplectic Hamiltonian formalism (MHF), whose main feature is to manifestly preserve the covariance of the Hamiltonian description of fields through the introduction of a generalized momentum per field derivative. The seminal papers on the MHF were published by de Donder [12] and by Weyl [13] in the 1930s. Nonetheless, after nearly a century of existence, the use of this formalism to construct classical or quantum field theories is not common practice, since its development beyond the basic definitions still presents formal lacunae [14]. In particular, an important drawback of the MHF is that it allows to define more than one Poisson bracket (PB) [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], which significantly dissipated the initial enthusiasm for using this formalism as the basis for developing a Lorentz-covariant canonical quantization.
Nevertheless, it has recently come to our attention that a Lorentz-covariant PB can be constructed in a straightforward manner for the electromagnetic radiation field using the MHF, by translating the multisymplectic description of the field into its momentum representation [25]. Additionally, and in a completely independent line of research, recent studies by Cetto et al. have established, within the framework of stochastic electrodynamics, a connection between the matrix formalism of quantum mechanics and the response functions of a particle interacting with an electromagnetic radiation field containing the zero-point radiation field, the non-thermal and stochastic electromagnetic background field that exists even in the absence of matter (for a review of stochastic electrodynamics up to 2015, see Ref. [26]). This connection is demonstrated by expressing the canonical variables of the mechanical subsystem in terms of the normal variables of the field’s modes of oscillation using PB. Thus, in the stationary phase of the interaction, where the field fully governs the particle’s motion, the kinematic transformation that changes the meaning of the quantities involved in the description of the particle becomes manifest [27, 28, 29]. Furthermore, the creation and annihilation operators of the radiation field have been deduced in similar manner, thereby providing a physical foundation for canonical quantization [30].
These results motivate us to revisit the problem of constructing a Lorentz-covariant canonical quantization process using the MHF, since a relativistic generalization of Cetto’s approach to quantization applicable to quantum field theories, most notably quantum electrodynamics, would not only physically justify the necessity for using operators instead of functions to describe fields, but also shed light on how to achieve a Lorentz-covariant canonical quantization. Hence, the purpose of the present work is to explore the possibility of describing a system composed of an electromagnetic radiation field and a point particle by means of a single Lorentz-covariant PB as a first step for developing the aforementioned generalization.
@@
The paper is organized as follows. In Sect. 2, after briefly reviewing the fundamentals of describing a system composed of an electromagnetic radiation field and a single point particle using the MHF, we translate the generated description to the momentum representation of the field. The central part of the paper is contained in Sect. 3 and deals with the proposal of a PB with respect to the canonical variables of the field that is capable not only of satisfying the restricted principle of relativity but also of reproducing well-established results from the conventional formulations of field theories. Subsequently, in Sect. 4, the non-covariant form of the proposal is analyzed and compared with the standard (non-covariant) PB that is associated with an electromagnetic radiation field. Lastly, we outline a description of the complete system on the basis of the found PB and briefly discuss its main implications in Sect. 5.
@@ 2 Multisymplectic Hamiltonian Description
We start by properly setting up the system to be investigated as an electromagnetic radiation field and a single point particle interacting with each other in a four-dimensional Minkowski space-time, with metric tensor $\eta_{\mu\nu}$ and signature (+,,,)(+,-,-,-). The action of such a system can be stated as
@@
where $\mathcal{L}_{\mathrm{field}}$, $\mathcal{L}_{\mathrm{part}}$, and $\mathcal{L}_{\mathrm{int}}$ are the Lagrangian densities associated with the free propagation of the field, the free motion of the particle, and the particle++field interaction, respectively. Thus, we establish that
$\mathcal{L}_{\mathrm{field}}=-\frac{1}{16\pi{c}}(F_{\mu\nu}F^{\mu\nu}+2\zeta(\partial_{\mu}A^{\mu})^{2}),$ (2)
$\mathcal{L}_{\mathrm{part}}=-\frac{m_{0}}{2}\int{d}\tau\hskip 1.42262pt\dot{u}_{\mu}\dot{u}^{\mu}\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (3)
@@
with $\zeta$ a positive, real parameter, $A_{\mu}$ the electromagnetic vector potential, $F_{\mu\nu}$ the electromagnetic tensor, and $m_{0}$, $e$, $u_{\mu}$, and $\tau$ the rest mass, electric charge, position, and proper time of the point particle, respectively. As usual, the movement of the particle is subject to the constraint $\dot{u}_{\mu}\dot{u}^{\mu}=c^{2}$. We have included a gauge-fixing term in the Lagrangian density $\mathcal{L}_{\mathrm{field}}$ in order to illustrate the descriptive capacity of the MHF
A canonical and manifestly Lorentz-covariant description of the system can then be achieved by associating the generalized (multisymplectic) momentum
$\theta_{\mu\nu}=-\frac{1}{4\pi{c}}(F_{\mu\nu}+\zeta\eta_{\mu\nu}\partial_{\alpha}A^{\alpha})$ (5)
to the electromagnetic vector potential and the well-known canonical momentum
$p_{\mu}=-m_{0}\dot{u}_{\mu}-\frac{e}{c}A_{\mu}(u_{\alpha})$ (6)
to the position of the particle, so that the action of the system can be written as
$\int{d}^{4}x\hskip 1.42262pt\left(\theta_{\mu\nu}\partial^{\mu}A^{\nu}+\int{d}\tau\hskip 1.42262ptp_{\mu}\dot{u}^{\mu}\delta^{(4)}(x_{\alpha}-u_{\alpha})-\mathcal{H}\right),$ (7)
where
$\mathcal{H}=-\frac{1}{2}\pi{c}\hskip 0.7113pt\theta_{\mu\nu}\left[\left(1+\frac{1}{2\zeta}\right)\theta^{\mu\nu}-\left(1-\frac{1}{2\zeta}\right)\theta^{\nu\mu}\right]-\frac{1}{2m_{0}}\int{d}\tau\left[p_{\mu}+\frac{eA_{\mu}}{c}\right]\left[p^{\mu}+\frac{eA^{\mu}}{c}\right]\delta^{(4)}(x_{\alpha}-u_{\alpha}),$ (8)
is the so-called de Donder–Weyl function, i.e., the Lorentz-covariant (multisymplectic) Hamiltonian density [12, 13]. On this basis, the Lorentz-covariant Hamilton’s equations
$\partial_{\mu}A_{\nu}=\frac{\partial\mathcal{H}}{\partial\theta^{\mu\nu}},\quad\partial^{\mu}\theta_{\mu\nu}=-\frac{\partial\mathcal{H}}{\partial{A}^{\nu}},$ (9)
and
$\int{d}\tau\hskip 1.42262pt\dot{u}_{\mu}=\frac{\partial}{\partial{p}^{\mu}}\int{d}^{4}x\hskip 1.42262pt\mathcal{H},\quad\int{d}\tau\hskip 1.42262pt\dot{p}_{\mu}=-\frac{\partial}{\partial{u}^{\mu}}\int{d}^{4}x\hskip 1.42262pt\mathcal{H},$ (10)
consistently reproduce the manifestly Lorentz-covariant forms of Maxwell’s equations and the equation of motion of the particle, respectively.
@@
$A_{\mu}=\frac{1}{8\pi^{3}}\lim_{m\rightarrow 0}\left[\delta_{\mu}^{\nu}+\frac{\zeta-1}{\zeta}\frac{d}{dm^{2}}\partial_{\mu}\partial^{\nu}\right]\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha}-m^{2}){\mathcal{A}}_{\nu}e^{-{i}k_{\alpha}x^{\alpha}}+\mathrm{c.c.},$ (11)
where the Dirac $\delta$-function $\delta(k_{\mu}k^{\mu})$ and the Heaviside step function $\Theta(k_{0})$ ensure that the dispersion relation $k_{0}=|\mathbf{k}|$ is satisfied, and ${\mathcal{A}}_{\mu}$ is a vector amplitude given by
${\mathcal{A}}_{\mu}={\mathcal{A}}^{\mathrm{free}}_{\mu}+4\pi{i}e\int{d}\tau\hskip 1.42262pt\dot{u}_{\mu}\Theta(x_{0}-u_{0})e^{ik_{\alpha}u^{\alpha}},$ (12)
with $\mathcal{A}_{\mu}^{\mathrm{free}}$ a vector amplitude depending solely on $k_{\mu}$. Having expressed the solution of the field equation in this manner, one can easily observe that the presence of the particle results for the field in a superposition of plane waves with constant amplitudes due to its free propagation and of waves with variable amplitudes whose origin can be shown to be the radiation reaction of the particle (for a detailed proof, see Barut [31]).
We aim to investigate the possibility of generalizing Cetto’s approach to quantization for the electromagnetic radiation field; hence, from this point on, we shall limit ourselves to working in the Lorentz gauge, i.e., taking $\zeta=1$. Then, recalling that the description of the field provided by the MHF in the momentum representation has been shown to be equivalent to its Lagrangian counterpart [25], we have that Eq. (8) can be rewritten as
@@
with $\delta_{k}$ a $\delta$-function such that $\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\delta_{k}=8\pi^{3}$, is the image of the system’s de Donder–Weyl function in the field’s momentum representation. This last function has the same information about the system as $\mathcal{H}$ and therefore constitutes the basis for this alternative description. Furthermore, the structure of Eq. (14) reveals that the de Donder–Weyl function is a Lorentz scalar proportional to the rest mass of the complete system, i.e., particle++field [^2] The origin of the proposed interpretation becomes evident when recalling that the dispersion relation for a massive field is $k_{\mu}k^{\mu}=m^{2}$, with $m$ being the mass parameter associated with the field. For purposes of comparison, examples of de Donder–Weyl functions in the momentum representation for other fields may be seen in Ref. [25]. Naturally, this identification implies a null value for the part of the de Donder–Weyl function that is associated with the free propagation of the electromagnetic radiation field since the dispersion relation states in such a case that $k_{\mu}k^{\mu}=k_{0}^{2}-|\mathbf{k}|^{2}=0$.
The image of the de Donder–Weyl function in the momentum representation can equivalently be expressed in terms of the field’s canonical variables ${q}_{\mu}$ and $\pi_{\mu\nu}$ [25], defined as
${q}_{\mu}=\frac{i(\mathcal{A}_{\mu}-\mathcal{A}_{\mu}^{\ast})}{\sqrt{8\pi{c}k_{0}}},\quad\pi_{\mu\nu}=\frac{k_{\mu}(\mathcal{A}_{\nu}+\mathcal{A}_{\nu}^{\ast})}{\sqrt{8\pi{c}k_{0}}},$ (15)
and whose Lorentz-covariant Hamilton’s equations are
@@
respectively. As a consequence, the set of equations constituted by Eqs. (10) and (16) completely governs the electrodynamics of the whole system in the Lorentz gauge.
@@ 3 Proposal for a Lorentz-covariant Poisson Bracket
In the Hamiltonian formulation of classical mechanics, the evolution equation of any phase-space function is stated in terms of its PB with the Hamiltonian function of the mechanical system concerned, with Hamilton’s equations being special cases of such an evolution equation for the canonical positions and momenta. In particular, when considering the Hamiltonian function itself, one immediately finds that its evolution equation reduces to an equality between its total and explicit time derivatives. So, if we claim that the above description of the system is canonical, we must verify whether the de Donder–Weyl function satisfies a similar mathematical relationship between its derivatives. To this end, we observe that it can be established straightforwardly from Eqs. (13) and (16) that
@@
so that we conclude that the de Donder–Weyl function indeed satisfies an equality such as the one mentioned.
The former statement brings PB back to the center of our discussion. As mentioned in Sect. 1, the MHF accepts multiple ways of defining such geometric objects, yet the structure of Eq. (17) strongly suggests proposing the bilinear form
$\{A,B\}=\int{d}^{4}k\hskip 1.42262ptV_{\mu}\left[\frac{\partial{A}}{\partial{q}^{\nu}}\frac{\partial{B}}{\partial\pi_{\mu\nu}}-\frac{\partial{B}}{\partial{q}^{\nu}}\frac{\partial{A}}{\partial\pi_{\mu\nu}}\right],$ (18)
@@
$\{q_{\mu}(x_{0},k_{\alpha}),\pi_{\lambda\nu}(x_{0},k_{\alpha}^{\prime})\}=V_{\lambda}\eta_{\mu\nu}\delta^{(4)}(k_{\alpha}-k_{\alpha}^{\prime}),$ (19)
@@
In order to get a hint on the structure of $V_{\mu}$, it is very enlightening to calculate the equal-time PB of the field and its (generalized) conjugate momentum,
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{1}{2}\left[\frac{\delta(0)}{8\pi^{3}}\right]\eta_{\mu\nu}K_{\lambda}(\mathbf{x},\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part},$ (20)
@@
$K_{\mu}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{4\pi^{3}}\int{d}^{4}k\hskip 1.42262pt\Theta(k_{0})\delta(k_{\alpha}k^{\alpha})\left[V_{\mu}\sin(\mathbf{k}\cdot\mathbf{x})\sin(\mathbf{k}\cdot\mathbf{x}^{\prime})+\frac{k_{\alpha}V^{\alpha}}{k_{\alpha}k^{\alpha}}k_{\mu}\cos(\mathbf{k}\cdot\mathbf{x})\cos(\mathbf{k}\cdot\mathbf{x}^{\prime})\right],$ (21)
since this result is independent of the auxiliary variables used to describe the system. Thus, Eq. (20) calls for taking $V_{\mu}$ equal to $ak_{\mu}$, with $a$ a proportionality constant, so that
$\{A_{\mu}(x_{0},\mathbf{x}),\theta_{\lambda\nu}(x_{0},\mathbf{x}^{\prime})\}=\frac{a}{2}\left[\frac{\delta(0)}{8\pi^{3}}\right]\eta_{0\lambda}\eta_{\mu\nu}\delta^{(3)}(\mathbf{x}-\mathbf{x}^{\prime})+\mathrm{non}\text{-}\mathrm{free}\hskip 1.42262pt\mathrm{part}.$ (22)
With this selection, we easily identify the well-known result from standard field theory for the mentioned PB in the part of the latter equation that is associated with the free propagation of the field.